OpenStudy (anonymous):

The rate law for a hypothetical reaction is rate = k [A]^2. When the concentration is 0.10 moles/liter, the rate is 2.7 × 10^-5 M*s^-1. What is the value of k?

OpenStudy (anonymous):

rate = k [A]^2 just plug in 2.7*10^(-5) = k*.1^2 ^rate ^.1 molles/liter maybe and then solve for k :)

OpenStudy (anonymous):

k = 0.27e-2

OpenStudy (anonymous):

umhm so wait... would it be the same as 2.7*10^-3M^-1s^-1?

OpenStudy (anonymous):

well unites should be M*s^-1 so ( moles/liter)^2* k's units= M*s^-1

OpenStudy (anonymous):

umhm alright. Thank you!

OpenStudy (anonymous):

is moles/liter=M?

OpenStudy (anonymous):

remember the unit M=mol/L \[ 2.7\times10^{-5}\,{\rm M/s}=k\times \left(0.1\,{\rm M}\right)^2\\ k=\frac{2.7\times10^{-5}}{0.01}\times \frac{{\rm M}}{\rm s\cdot M^2}\\ \boxed{k=2.7\times10^{-3}{\rm M^{-1}\cdot s^{-1}}} \] the reaction is thus, a second order reaction.

OpenStudy (anonymous):

Soo, I was right? :D woho! Thanks alot electrokid!

OpenStudy (anonymous):

yaya. good job.